OpenStudy (amonoconnor):

A person weighing 800 Newtons travels to another planet with twice the mass and twice the radius of Earth. What is the person's weight on the new plant?

OpenStudy (anonymous):

400N

OpenStudy (amonoconnor):

Can you explain how you got there a bit? What equation/ formula did you use?

OpenStudy (amonoconnor):

@MayankD :)

OpenStudy (anonymous):

The weight of a body is basically the gravitational force acting on it, Newtons law of gravitation states that force of attraction between two particles of mass M and m is given by F=-GMm/(R^2){- sign just denotes that force is attractive} where R is the distance between the two particles and G is the universal Gravitation constant Now in this case let the mass of the person be m. it remains constant regardless of which planet the person is on Let M be mass of Earth and R be radius of Earth[which is also distance between centre of eartha nd the person] so, 800=-GMm/(R^2)-------(1) Now the mass of the other planet is 2M and radius is 2R So, Weight=-G(2M)m/(2R)^2 Weight= Weight on earth/2=800/2=400N

OpenStudy (amonoconnor):

Alrighty, thank you so much!